### v1

# Advancing Physics Data Analysis through Machine Learning and Physics-Informed Neural Networks

Abstract
In an era increasingly focused on green computing and explainable AI, revisiting traditional approaches in theoretical and phenomenological particle physics is paramount. This project evaluates various machine learning (ML) algorithms—including Nearest Neighbors, Decision Trees, Random Forest, AdaBoost, Naive Bayes, Quadratic Discriminant Analysis (QDA), and XGBoost—alongside standard neural networks and a novel Physics-Informed Neural Network (PINN) for physics data analysis. We apply these techniques to a binary classification task that distinguishes the experimental viability of simulated scenarios based on Higgs observables and essential parameters. Through this comprehensive analysis, we aim to showcase the capabilities and computational efficiency of each model in binary classification tasks, thereby contributing to the ongoing discourse on integrating ML and Deep Neural Networks (DNNs) into physics research. In this study, XGBoost emerged as the preferred choice among the evaluated machine learning algorithms for its speed and effectiveness, especially in the initial stages of computation with limited datasets. However, while standard Neural Networks and Physics-Informed Neural Networks (PINNs) demonstrated superior performance in terms of accuracy and adherence to physical laws, they require more computational time. These findings underscore the trade-offs between computational efficiency and model sophistication.
1 Introduction
Navigating the enigmatic realms of particle physics often leads us to confront phenomena that remain shrouded in mystery, such as the puzzling properties of neutrino masses and the elusive essence of dark matter. These scientific enigmas challenge our current frameworks of understanding and compel the pursuit of more sophisticated theories. To unravel these mysteries, it is essential to test a variety of theoretical and phenomenological models against the experimental data. Our study is propelled to help in the phenomenological study of those models, by implementing a machine learning filter. Our approach harnesses the power of machine learning to refine the investigative lens through which we examine these models.
In the initial stage of the physics analysis, we begin by randomly initializing the free parameters of our theoretical model. These parameters undergo a preliminary process where the tree-level masses are computed, a crucial step for establishing the foundation of our simulation environment. Subsequently, these calculated values are inputted into the SPheno [1, 2] code. SPheno plays a pivotal role in our analysis, as it is tasked with a comprehensive calculation suite: it determines two- and three-body decay modes of both supersymmetric particles and Higgs bosons. Beyond decay calculations, SPheno extends its functionality to evaluate production cross-sections for these particles in electron-positron annihilation scenarios. The versatility of SPheno’s output allows for further analytical exploration through integration with other sophisticated software tools. Notably, it can interface with HiggsBounds [3] and HiggsSignals [4], which scrutinize the compatibility of Higgs boson observations with theoretical predictions, and with MadGraph5_aMC@NLO [5], a tool renowned for simulating high-energy collisions that further elucidates the dynamical behaviours of particles under our model. This methodical and integrated approach, leveraging the strengths of multiple computational tools, furnishes a comprehensive examination of the theoretical model, enabling a detailed exploration of its phenomenological implications which is seen in the flowchart in Fig. 1. The flowchart outlined above enables highly accurate calculations, but this precision comes with a significant time cost. To address this efficiency challenge, we propose incorporating a machine learning (ML) filter at the initial stage of variable initialization. This ML filter is designed to provide a preliminary assessment of whether the randomly initialized variables are likely to be experimentally viable. By applying this filter, we can selectively perform the more time-intensive, precise calculations only on those sets of parameters predicted to be within experimental bounds. This approach aims to streamline the computational process, ensuring that resources are focused on scenarios yielding meaningful results. While the detailed exploration of physics scenarios and the underlying calculations form the cornerstone of our work, they extend beyond the primary focus of this project. As such, we do not delve deeply into the specifics of these physics computations within the current study. For readers interested in a more thorough understanding of these aspects, we direct them to Ferreira et al. [6], where these topics are extensively covered. Our emphasis here is on leveraging machine learning techniques to enhance the efficiency and effectiveness of identifying experimentally viable parameter sets, rather than on the intricacies of the physics calculations themselves.
Initialization of VariablesPre-process of DataSPhenoMadGraph5_aMC@NLOHiggsBounds/HiggsSignals
Figure 1: Process Flowchart
2 Dataset Description and Preparation
The dataset employed in this study was derived from the comprehensive phenomenological analysis conducted by Ferreira et al. [6, 7], focusing on a Branco-Grimus-Lavoura (BGL) model. This model, a variant within the two Higgs doublet framework augmented by a single scalar singlet, stands out for its rich phenomenology in particle physics. The original dataset encompasses both experimental outcomes and the corresponding set of free parameters instrumental in computing these values.
For our analysis, we engaged in an extensive data-cleaning process. This refinement involved selecting only the vital free parameters and specific physics values (denoted by $\mathbf{P}$) relevant for predictive modelling. Additionally, the dataset was labelled based on a spectrum of experimental observables, categorizing each instance as either experimentally viable or nonviable according to predefined experimental criteria. This preprocessing was imperative for ensuring that our models concentrated on the free parameters, thus enhancing the accuracy and interpretability of the analysis within the framework of the BGL model. Notably, our dataset contains approximately 16% of viable points, indicating an imbalance. We represent the dataset as $\mathbf{X}$, where $\mathbf{X}=\{x_{1},\dots,x_{N}\}$, and N is the number of simulated data points. Correspondingly, a supplementary dataset contains the physics observables, represented as $\mathbf{P}$, such that for each data point $x_{i}$ with $i\in\{1,\dots,N\}$, there exists an associated set of physics observables $\mathbf{p}_{i}=\{p^{i}_{1},p^{i}_{2},p^{i}_{3}\}$, with $i$ ranging from 1 to N.
3 Methodology
As part of our study, we conducted a comprehensive evaluation of various machine learning algorithms to determine their effectiveness in predicting outcomes based on our dataset. The assessment process began with the preparation of the dataset, followed by the application of standard scaling to normalize the feature space. This preprocessing step is crucial for models that are sensitive to feature magnitude. We implemented a stratified k-fold cross-validation approach with five folds for evaluating the ML algorithms. This method ensures a robust assessment while maintaining the original dataset’s class proportion within each fold, thus addressing potential biases in model performance due to imbalanced class distributions. Our selection of ML models ranged from simpler algorithms, such as Gaussian Naive Bayes and Decision Trees, to more complex ensemble methods including Random Forest, AdaBoost, and XGBoost [8]. Each model was rigorously tested to gain a comprehensive understanding of its performance, focusing on metrics such as accuracy, precision, recall, F1 score, and the area under the ROC curve (ROC AUC). These metrics were selected for their capacity to provide a detailed insight into model performance, particularly in binary classification tasks where balancing true positive and false positive rates is essential. We systematically recorded the evaluation results, noting not only the mean and standard deviation of each metric across the folds but also the time required for each model’s training and evaluation. This method allowed us to make nuanced comparisons between the models, considering both their predictive performance and computational efficiency.
For the deep neural networks segment of our study, we adopted a focused approach to assess the training procedure, primarily utilizing the validation loss as a key indicator of performance. This method involves dividing our dataset into training, validation, and test subsets, with the latter serving to further validate our results and as a proxy to estimate the model’s generalization ability on unseen data. Throughout the training epochs, we closely monitored the validation loss, aiming to minimize it as an objective measure of the model’s prediction error. A distinctive emphasis was placed on exploring the capabilities of Physics-Informed Neural Networks (PINNs) [9] alongside traditional Neural Networks (NNs) to address the unique challenges presented by our physics-based dataset. In subsection 3.1, we delve into the structural details of these models and their operational dynamics.
Input LayerHidden Layer 1Hidden Layer 2Hidden Layer 3Output Layer
Input LayerHidden Layer 1Hidden Layer 2Physics-Informed LayerOutput Layer
Figure 2: Comparison of a standard Neural Network Fig 2(a) and a Physics-Informed Neural Network (PINN) Fig 2(b).
3.1 Physics-Informed Neural Networks (PINNs)
The cornerstone of Physics-Informed Neural Networks (PINNs) lies in their innovative integration of physical laws and information directly into the learning process. This integration is meticulously encapsulated within our dual-output model architecture. This architecture, visualized in Fig 2, encapsulates the essence of PINNs and illustrates the differences with traditional neural networks (NNs). The PINN enables the simultaneous prediction of physics-informed outputs $\hat{\mathbf{P}}$ alongside the primary task-specific predictions $\hat{\mathbf{Y}}$, ensuring compliance with underlying physical principles:
$\begin{split}\hat{\mathbf{P}}&=\mathcal{G}_{w}(\mathbf{X})\Leftrightarrow\hat{\mathbf{p}}_{i}=\mathcal{G}_{w}(x_{i}),\\ \hat{\mathbf{Y}}&=\mathcal{F}_{w}(\mathbf{X},\hat{\mathbf{P}})\Leftrightarrow\hat{y}_{i}=\mathcal{F}_{w}(x_{i},\hat{\mathbf{p}}_{i}).\end{split}$ (1)
The critical innovation in our PINN approach (Fig 3) occurs during the training phase (Alg 1), particularly in the structuring of the loss function. In addition to the conventional loss term for measuring the accuracy of primary task predictions (e.g., classification accuracy or mean squared error for regression), an additional term is introduced to ensure that the predictions made by the physics-informed layer—P1, P2, P3 as shown in Fig 3—closely match the actual physics quantities derived from the data or known physical laws. To effectively balance the contribution of this physics-based loss term with the primary task loss, we include a normalization constant, $\lambda$. In this study, $\lambda$ is set to one, simplifying the integration process of these terms and maintaining a direct influence of the physics-informed constraints on the overall loss function.
$\begin{split}\mathcal{L}_{\text{total}}&=\mathcal{L}_{\text{classification}}+\lambda\mathcal{L}_{\text{physics}},\\ &=\mathcal{L}_{\text{classification}}(\mathbf{Y},\hat{\mathbf{Y}})+\lambda\mathcal{L}_{\text{physics}}(\mathbf{P},\hat{\mathbf{P}}),\\ &=\frac{1}{N}\sum_{i}^{N}{\mathcal{L}_{\text{classification}}(y_{i},\mathcal{F}_{w}(x_{i},\mathcal{G}_{w}(x_{i})))+\lambda\mathcal{L}_{\text{physics}}(\mathbf{p}_{i},\mathcal{G}_{w}(x_{i}))}.\end{split}$ (2)
By incorporating this additional physics-based loss term, the PINN is explicitly trained not only to achieve accuracy in its primary tasks but also to approximate physical values and make decisions informed by physics. This approach enhances both the interpretability and reliability of the model’s outputs, bridging the gap between complex data analysis and fundamental physical principles.
The training process diverges from the evaluation process, by focusing exclusively on the classification loss during the evaluation phase. This strategic omission of the physics loss term during model evaluation allows for a direct and clear comparison between the performance metrics of standard Neural Networks (NNs) and Physics-Informed Neural Networks (PINNs), ensuring our evaluations are objective and insightful.
InputH1-1H1-2H1-NH2-1H2-2H2-NP1P2P3Output
Figure 3: Architecture of the Physics-Informed Neural Network (PINN). The network comprises an input layer, followed by multiple fully connected hidden layers (H1 and H2), and a specialized physics-informed layer (highlighted in red), concluding with an output layer. While the illustration shows selective connections for clarity, it is important to note that each layer is fully connected to its subsequent layer. The physics-informed layer (P1, P2, P3) is designed to encode specific physical principles or constraints relevant to the problem domain, facilitating the integration of domain knowledge directly into the learning process.
Data: Training data $\mathbf{X}=\{x_{1},\dots,x_{N}\}$, Training labels $\mathbf{Y}=\{y_{1},\dots,y_{N}\}$, Physics data $\mathbf{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{N}\}$
Training: Tuning of $w$ based on $\hat{y}_{i}=f_{w}(x_{i},p_{i})$ and $\hat{\mathbf{p}}_{i}=\mathcal{G}_{w}(x_{i})$;
// Start of training step
for all $x_{i}\ \in\ X$ with $i\ \in\{1,\dots,N\}$ do
      Forward pass to compute predictions $\hat{y}_{i}$ and physics-informed outputs $\hat{p}_{i}$;
      Compute classification loss $\mathcal{L}_{\text{classification}}(y_{i},\hat{y}_{i})=\mathcal{L}_{\text{classification}}(y_{i},f_{w}(x_{i},\hat{p}_{i}))$;
      Compute physics loss $\mathcal{L}_{\text{physics}}(p_{i},\hat{\mathbf{p}}_{i})=\mathcal{L}_{\text{physics}}(p_{i},\mathcal{G}_{w}(x_{i}))$;
      Compute total loss $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{classification}}+\lambda\mathcal{L}_{\text{physics}}$;
      Backpropagation;
end
Algorithm 1 Training a Physics-Informed Neural Network (PINN)
4 Results and Discussion
Model Accuracy Precision Recall F1 Score ROC AUC Time (seconds)
Nearest Neighbors 0.8359 $\pm$ 0.0026 0.5160 $\pm$ 0.0091 0.4461 $\pm$ 0.0074 0.4785 $\pm$ 0.0033 0.7924 $\pm$ 0.0066 629.70
Decision Tree 0.8781 $\pm$ 0.0007 0.6851 $\pm$ 0.0039 0.5140 $\pm$ 0.0060 0.5874 $\pm$ 0.0037 0.9015 $\pm$ 0.0032 149.38
Random Forest 0.8333 $\pm$ 0.0063 0.3493 $\pm$ 0.8596 0.0130 $\pm$ 0.0401 0.0249 $\pm$ 0.0761 0.8463 $\pm$ 0.0090 27.13
AdaBoost 0.8576 $\pm$ 0.0026 0.6488 $\pm$ 0.0243 0.3409 $\pm$ 0.0270 0.4466 $\pm$ 0.0202 0.8804 $\pm$ 0.0021 580.41
Naive Bayes 0.6059 $\pm$ 0.0045 0.2953 $\pm$ 0.0024 0.9636 $\pm$ 0.0042 0.4521 $\pm$ 0.0029 0.8174 $\pm$ 0.0048 2.01
QDA 0.6183 $\pm$ 0.0058 0.3012 $\pm$ 0.0034 0.9562 $\pm$ 0.0100 0.4581 $\pm$ 0.0045 0.8173 $\pm$ 0.0034 8.01
XGBoost 0.9608 $\pm$ 0.0014 0.8990 $\pm$ 0.0121 0.8651 $\pm$ 0.0105 0.8817 $\pm$ 0.0037 0.9882 $\pm$ 0.0009 9.24
Table 1: Performance Metrics of Various Machine Learning Models in Physics Data Analysis. This table presents a comparative overview of the accuracy, precision, recall, F1 score, ROC AUC, and computation time for each model. Top-performing values in each category are emphasized in bold.
In our evaluation of machine learning models for physics data analysis which can be seen in Table 1, we focused on critical metrics such as Precision, Recall, and F1 Score, which are essential for accurately identifying true positive scenarios. XGBoost excelled among the models tested, achieving the highest F1 Score (0.8817 ± 0.0037). This score demonstrates an optimal balance between precision and recall, pivotal for minimizing classification errors in our study. Moreover, XGBoost also led with the best ROC AUC score (0.9882 ± 0.0009), underscoring its exceptional ability to differentiate between classes across various threshold settings. It additionally recorded the highest precision (0.8990 ± 0.0121). While the Naive Bayes model displayed the highest recall, it did so at the expense of precision, highlighting the trade-offs involved in model selection. Overall, XGBoost proved to be the most effective model for our specific analytical requirements, offering high reliability, robust discrimination capabilities, and an outstanding balance of performance metrics, requiring also 9.24 sec for training.
Learning Rate Accuracy Precision Recall F1 Score ROC AUC
0.0001 0.9740 0.9404 0.9040 0.9219 0.9938
0.0002 0.9736 0.9531 0.8879 0.9193 0.9936
0.0005 0.9716 0.9217 0.9099 0.9158 0.9937
8e-05 0.9682 0.9549 0.8530 0.9011 0.9924
Table 2: Performance Metrics of Top Neural Network Models Under Different Learning Rates. This table showcases the results for neural network configurations that demonstrated the best overall performance in our analysis. Values in bold represent the highest scores achieved in each respective performance category across all tested models, indicating the most effective model settings for each metric."
Our analysis of the neural network models which can be seen in Table 2, examined under various learning rates, reveal small yet crucial differences in performance metrics that impact their efficacy in scenario classification. The model with a learning rate of 0.0001 outperforms others, achieving the highest F1 Score (0.9219) and Precision (0.9404), alongside an impressive Recall (0.9040), indicating its strong capability to balance between identifying true positives and minimizing false outcomes. Its ROC AUC of 0.9938 also demonstrates superior discriminatory power.
In our study, while XGBoost offers a balanced performance across all metrics, it is outpaced by almost all the neural network models tested. Notably, the neural network with a learning rate of 0.0001 not only demonstrated superior performance with the highest F1 Score and Precision but also achieved an impressive Recall (0.9040), indicating its exceptional capability in balancing true positive identification and minimizing false outcomes. Its ROC AUC of 0.9938 further underscores its superior discriminatory power over XGBoost. Although these neural networks generally require more training time than XGBoost, their superior performance across almost all metrics makes them highly effective for applications where model accuracy is prioritized over training speed.
LR Act. Func. Phys. Act. Func. Accuracy Precision Recall F1 Score ROC AUC
0.0009 GELU GELU 0.9711 0.9145 0.9152 0.9148 0.9933
0.004 GELU GELU 0.9684 0.8853 0.9351 0.9095 0.9924
0.005 ReLU LeakyReLU 0.9694 0.9178 0.9003 0.9089 0.9894
0.004 ReLU LeakyReLU 0.9706 0.9013 0.9286 0.9147 0.9924
0.006 ReLU ELU 0.9619 0.8546 0.9345 0.8928 0.9913
0.003 ReLU ELU 0.9608 0.8705 0.9032 0.8866 0.9910
0.005 GELU GELU 0.9630 0.8605 0.9331 0.8953 0.9904
0.003 GELU GELU 0.9658 0.8746 0.9319 0.9024 0.9926
0.002 GELU GELU 0.9675 0.8894 0.9235 0.9061 0.9899
0.006 ReLU LeakyReLU 0.9676 0.9172 0.8894 0.9031 0.9911
Table 3: Performance Comparison of PINN Models with Varying Activation Functions. This table displays the accuracy, precision, recall, F1 scores, and ROC AUC for Physics-Informed Neural Networks (PINNs) across different learning rates and activation function configurations. The best values in each performance category are highlighted in bold, emphasizing the models’ top performances in our evaluation.
In our expanded analysis of Physics-Informed Neural Network (PINN) models detailed in Table 3, we assessed various configurations to gauge how learning rates and activation functions impact essential metrics like Precision, Recall, and the F1 Score. The top PINN model, with a learning rate of 0.0009 and GELU as both the activation and physics activation function, stands out by achieving the highest F1 Score (0.9148). This score reflects an optimal balance, with equally high Precision (0.9145) and Recall (0.9152), indicating effective integration of physical principles into the learning process. The superior performance of GELU can be attributed to its Gaussian Distribution-Based Smoothing, which weights inputs through a probabilistic lens, enhancing the model’s ability to handle the complex, non-linear relationships inherent in physics data. Additionally, GELU supports smoother gradient flow during backpropagation compared to other activation functions. This characteristic prevents issues related to vanishing or exploding gradients, facilitating more stable and effective learning across deeper network architectures. These properties make GELU particularly suitable for PINNs, where capturing subtle physical behaviors accurately is crucial.
Comparatively, this top PINN model outperforms the best XGBoost configuration, which had an F1 Score of 0.8817, demonstrating superior balance and precision-recall trade-off. When matched against the top-performing neural network model, which reached an F1 Score of 0.9219, the leading PINN model shows competitive precision but a slightly lower overall F1 score. Despite this, the unique strength of PINNs in incorporating physical laws directly into their framework provides an explainability advantage against both XGBoost and top-performing neural network.
These results highlight the effectiveness of PINNs in optimizing both sensitivity and specificity, crucial for reducing false negatives and positives in complex systems and the ability to make decisions based on physics information without losing effectiveness. Such capabilities make PINNs particularly valuable in physics data analysis, where ensuring high accuracy in prediction under physical laws is paramount.
In summary, for the initial stages of computation, where only a small portion of the dataset is available, XGBoost is highly recommended as the machine learning filter. Its speed and relative accuracy make it ideal for quick, preliminary filter. On the other hand, for long-term simulations where comprehensive datasets are collected, Physics-Informed Neural Networks (PINNs) are advisable. PINNs leverage the underlying physical laws to enhance the predictive capability of traditional neural networks, integrating crucial physics-based insights into the otherwise non-transparent ML models. This approach ensures that the models are not only accurate but also fulfil the established physical principles, making them particularly valuable for in-depth, physics-intensive investigations.
5 Future Work
Looking ahead, the exploration of several intriguing research directions promises to enrich the domain of physics data analysis further. One of the most compelling prospects involves investigating the transferability of the top-performing neural network and PINN models to other physical systems with similar Lagrangians. This endeavour would not only highlight the generalizability and adaptability of these models but also deepen our comprehension of their fundamental workings. Moreover, determining the minimal dataset size necessary to achieve approximately 90% accuracy emerges as a fascinating challenge, especially in the case of XGBoost, which we recommended as a preliminary filter. This inquiry would entail a comparison between the time savings from reduced simulation requirements and the computational demands of model training, aiming to optimize data usage efficiency. Additionally, the integration of uncertainty quantification within PINNs could significantly enhance their predictive power by offering probabilistic bounds on predictions, thereby bolstering confidence in model outputs and facilitating more informed decision-making. The adoption of ensemble methods that fuse the strengths of diverse models, including PINNs and traditional machine learning algorithms, may also unveil synergistic benefits, potentially elevating overall model performance and robustness. Lastly, applying PINNs in applications, spanning from particle physics experiments to cosmological data analysis, would furnish a comprehensive testing ground for these advanced computational tools, pushing the frontiers of current technology and knowledge in a concerted effort to unravel the complexities of our universe.